OpenStudy (anonymous):

Let theta be an angle in quadrant III such that cos theta=-3/5 . Find the exact values of csc theta and tan theta .

OpenStudy (amistre64):

3-4-5- lol

OpenStudy (anonymous):

draw a triangle. other side will be 4 by pythagoras so you have a 3 - 4 - 5 right triangle

OpenStudy (anonymous):

quadrant III (i love the roman numerals) so sine will be negative as is cosine

OpenStudy (anonymous):

\[\cos(\theta)=-\frac{3}{5}\] so \[\sin(\theta)=-\frac{4}{5}\]

OpenStudy (anonymous):

cos theta =-3/5 so theta=126.86 csc 126.86=1.24 tan126.86=-1.334

OpenStudy (anonymous):

\[\tan(\theta)=\frac{4}{3}\] the minus signs and the denominators cancel

OpenStudy (anonymous):

\[\csc(\theta)=\frac{1}{\sin(\theta)}=-\frac{5}{4}\]

OpenStudy (anonymous):

which, btw, is not 1.24 but rather 1.25 so forget about finding theta and using a calculator, just find the ratios

OpenStudy (anonymous):

and in any case it is negative. you are in quadrant III so there is no way theta is 126.86. this is the danger of using a calculator with inverse trig functions so don't do it.

OpenStudy (anonymous):

oh i guess that was all. but it should be clear than once you have cosine and know the quadrant you can find all the other trig functions yes?